OpenStudy (anonymous):

1. The Mathematics Department at Sun Tan University is sending a delegation of mathematicians and graduate students to a conference. They will pay for 4 mathematicians and 5 graduate students to go. If the department has 16 mathematicians and 24 graduate students, how many different delegations are possible?

OpenStudy (nurali):

The delegation consists of a combination of 4 mathematicians and 5 graduate students. 16C4 * 24C5 = [16! / (12! 4!)] [24! / (19! 5!)] = 16! 24! / (12! 4! 19! 5!) = 16*15*14*13 * 24*23*22*21*20 / (24 * 120) = 23*22*21*20 * 16*15*14*13 / 120 = 77,357,280 delegations possible

OpenStudy (nurali):

Welcome to Openstudy.